# Extremely Amenable Automorphism Groups of Countable Structures

### v1

@@ 6. Some Descriptive Set Theoretic Results
For any $A\subseteq\mathbb{R}^{\mathbb{N}}$, let $\sim_{A}$ denote $\sim\upharpoonright(A\times A)$, and similarly, let $\approx_{A}$ denote $\approx\upharpoonright(A\times A)$.
Thus $\approx_{\mathsf{CDVS}}$ is the equivalence between countable distance value sets and $\approx_{\mathsf{CDVS}_{0}}$ is the equivalence between countable distance values sets without a positive lower bound. Theorem 4.5 can be restated as $\sim_{\mathsf{CDVS}_{0}}\,=\,\approx_{\mathsf{CDVS}_{0}}$.
The following theorem is a summary of our results about these equivalence relations.
Theorem 6.2.
$E_{GL_{2}(\mathbb{Q})}^{IR}\,\leq_{B}\ \sim_{\mathsf{CDVS}_{0}}\ \leq_{B}\ \approx_{\mathsf{CDVS}}\ \leq_{B}\ \cong_{\mathsf{EA}}$.
Proof.
$E_{GL_{2}(\mathbb{Q})}^{IR}\leq_{B}\,\sim_{\mathsf{CDVS}_{0}}$ is witnessed by the map $\alpha\mapsto\Delta_{\alpha}$ and guaranteed by Theorem 4.6. We only need to see that there is a canonical Borel function that gives a coding sequence for $\Delta_{\alpha}=\{p\alpha+q\colon p,q\in\mathbb{Q}\}\cap(0,+\infty)$. One way to do this is as follows. Set the odd part of the sequence to zero. For the even part of the sequence, go through a canonical enumeration of all pairs $(p,q)$ where $p,q\in\mathbb{Q}$, and include $p\alpha+q$ only when it is a positive number.
$\sim_{\mathsf{CDVS}_{0}}\,\leq_{B}\,\approx_{\mathsf{CDVS}}$ is clear in view of Theorem 4.5, which states $\sim_{\mathsf{CDVS}_{0}}\,=\,\approx_{\mathsf{CDVS}_{0}}$, and the obvious $\approx_{\mathsf{CDVS}_{0}}\,\leq_{B}\,\approx_{\mathsf{CDVS}}$.
The last reduction $\approx_{\mathsf{CDVS}}\,\leq_{B}\,\cong_{\mathsf{EA}}$ is witnessed by the map $\Delta\mapsto\mbox{\rm Aut}(\mathbb{U}^{<}_{\Delta})$ and guaranteed by Theorem 5.4. Again we only need to show that the map from a distance value set $\Delta$ to ${\rm Aut}(\mathbb{U}_{\Delta}^{<})$ is Borel. This is done in two steps. In the first step, we describe a Borel way to enumerate the class $\mathcal{K}^{<}_{\Delta}$, the class of all finite ordered $\Delta$-metric spaces, from a given sequence $(d_{i})_{i}$ coding $\Delta$. For this, we fix a countable language $\mathcal{L}$ consisting of countably many binary relation symbols, i.e., =$\mathcal{L}=\{R_{i}\}_{i\in\mathbb{N}}\cup\{<\}$. Enumerate those finite $\mathcal{L}$-structures $\mathcal{A}$ such that
  1. (i)
  2. (ii)
  3. (iii)
    $<^{\mathcal{A}}$ is a linear order.
Now, given an enumeration of $\mathcal{K}^{<}_{\Delta}$, one can construct the Fraïssé limit in a Borel way. This follows from the existence proof of Fraïssé’s theorem (see, e.g. Lemma 7.1.6 of Hodges [13]). Since the Fraïssé structure is ultrahomogeneous, its automorphism group intersects a basic open set $N_{\bar{a},\bar{b}}$ iff the substructures $\bar{a}$ and $\bar{b}$ are isomorphic. Thus the map from the Fraïssé structures to their automorphism groups is Borel. ∎
Finally, we compare the above equivalence relations with some benchmark equivalence relations in the theory of Borel reducibility for equivalence relations.
We list all relevant benchmark equivalence relations in the following. Discussions about the ones without references can be found in [7].
  1. (1)
    $=$ is the equality relation on $2^{\mathbb{N}}$.
  2. (2)
    $E_{0}$ is the eventual agreement relation on $2^{\mathbb{N}}$ defined as
    $(x_{i})_{i}E_{0}(y_{i})_{i}\iff\exists i\ \forall j\geq i\ x_{j}=y_{j}.$
  3. (3)
    $E_{\infty}$ is the universal countable Borel equivalence relation. It can be defined as the following equivalence relation on $2^{\mathbb{F}_{2}}$, where $\mathbb{F}_{2}$ is the free group with two generators:
    $(x_{g})_{g}E_{\infty}(y_{g})_{g}\iff\exists h\in\mathbb{F}_{2}\ \forall g\in\mathbb{F}_{2}\ x_{g}=y_{gh}.$
  4. (4)
    $=^{+}$ is the equivalence relation on $(2^{\mathbb{N}})^{\mathbb{N}}$ defined as
    $(x_{i})_{i}\ =^{+}(y_{i})_{i}\iff\{x_{i}\colon i\in\mathbb{N}\}=\{y_{i}\colon i\in\mathbb{N}\}.$
  5. (5)
    $=^{++}$ is the equivalence relation on $((2^{\mathbb{N}})^{\mathbb{N}})^{\mathbb{N}}$ defined as
    $(x_{i})_{i}\ =^{++}(y_{i})_{i}\iff\forall i\ \exists j\ x_{i}=^{+}y_{j}\mbox{ and }\forall j\ \exists i\ x_{i}=^{+}y_{j}.$
  6. (6)
    1. a)
      $A$ is a hereditarily countable set in $\mathcal{P}^{3}(\mathbb{N})$;
    2. b)
      • for any $a,b\in A$ there is some $r$ such that $R(a,b,r)$ holds;
    The equivalence relation $\cong^{*}_{3,k}$ is defined as the isomorphism relation of countable structures coding pairs $(A,R)$ in $Q^{*}_{k}(\mathbb{N})$. These equivalence relations were defined by Hjorth–Kechris–Louveau [12]. Our presentation follows [2].
The following Borel reducibility results are well-known (see [7]):
$=\ <_{B}E_{0}\,<_{B}E_{\infty}<_{B}\ =^{+}\,<_{B}\ =^{++}.$
It was also shown in Hjorth–Kechris–Louveau [12] that
$=^{+}\,<_{B}\ \cong^{*}_{3,0}\ <_{B}\ \cong^{*}_{3,1}\ <_{B}\ =^{++}.$
Corollary 5.5 can be restated as the equivalence relation $=$ being a lower bound for $\cong_{\mathsf{EA}}$ in the Borel reducibility hierarchy.
We note the following fact.
Lemma 6.3.
$E_{0}\leq_{B}E_{GL_{2}(\mathbb{Q})}^{IR}\leq_{B}E_{\infty}$.
Proof.
$E_{GL_{2}(\mathbb{Q})}^{IR}$ is induced by a continuous action of a countable group on a standard Borel space. Thus $E_{GL_{2}(\mathbb{Q})}^{IR}\leq_{B}E_{\infty}$. Note that every orbit is dense for this action. It follows from a theorem of Becker–Kechris (see Theorem 6.2.1 of [7]) that $E_{0}\leq_{B}E_{GL_{2}(\mathbb{Q})}^{IR}$. ∎
This gives $E_{0}$ as a lower bound for $\cong_{\mathsf{EA}}$, which is an improvement of Corollary 5.5.
@@
Next we turn to equivalence relations $\sim_{\mathsf{CDVS}_{0}}$. Here we consider another equivalence relation defined and studied by Calderoni–Marker–Motto Ros–Shani [2]. Let $\mathsf{ArGp}$ be the space of all countable Archimedean ordered groups, and let $\cong_{\mathsf{ArGp}}$ denote the isomorphism relation on $\mathsf{ArGp}$. It was proved in [2] that
$=^{+}\ <_{B}\ \cong_{\mathsf{ArGp}}\ \leq_{B}\ \cong^{*}_{3,1}.$
@@
To see that $\Delta$ is a Borel reduction witnessing $\cong_{\mathsf{ArGp}}\ \leq_{B}\ \sim_{\mathsf{CDVS}_{0}}$, we use a result of Hion (see Lemma 2.2 of [2]) which states that for any $G,H\in\mathsf{ArGp}$, $G\cong H$ iff there is a real number $\alpha$ such that $\alpha\cdot G=H$. It is then clear that $G\cong H$ iff there is a positive real number such that $\alpha\cdot G=H$ iff there is a positive real number such that $\alpha\cdot\Delta(G)=\Delta(H)$ iff $\Delta(G)\sim\Delta(H)$. ∎
Combining with the results of [2], we obtain $=^{+}$ as a non-optimal lower bound for $\sim_{\mathsf{CDVS}_{0}}$, as well as for $\cong_{\mathsf{EA}}$, in the Borel reducibility hierarchy. In symbols, we have $=^{+}\ <_{B}\ \cong_{\mathsf{EA}}$. This is a further improvement of Corollary 5.5 and Lemma 6.3. In particular, we have shown Theorem 1.7.
The next result will give an upper bound for $\sim_{\mathsf{CDVS}_{0}}$, and in particular implies
$=^{+}\ <_{B}\ \sim_{\mathsf{CDVS}_{0}}\ <_{B}\ =^{++}.$
Proposition 6.5.
@@
=$\mathcal{L}=\{0,c,\leq,+\}\cup\{R_{q}\colon q\in\mathbb{Q}^{+}\}$
@@
  • $+$ is a binary function symbol;
@@
In a countable distance value set $\Delta$ with $\inf\Delta=0$, we interpret $c$ to be 00 if $\sup\Delta=+\infty$ and to be $\sup\Delta$ otherwise. For each $q\in\mathbb{Q}^{+}$, $R_{q}$ is interpreted as $R_{q}(x,y)$ iff $q<x/y$. Under these interpretations, $\Delta\cup\{0\}$ becomes a countable $\mathcal{L}$-structure.
@@
    $\leq$ is a linear ordering of $|\mathcal{M}|$, 00 is the $\leq$-least element, and for all $x,x^{\prime}\in|\mathcal{M}|\setminus\{0\}$, $x\leq x^{\prime}$ iff $x=x^{\prime}$ or $R_{1}(x^{\prime},x)$.
@@
    For all $x,x^{\prime},y\in|\mathcal{M}|\setminus\{0\}$ and $q\in\mathbb{Q}^{+}$, $R_{q}(x+x^{\prime},y)$ iff there are $q_{1},q_{2}\in\mathbb{Q}^{+}$ with $q_{1}+q_{2}=q$ such that $R_{q_{1}}(x,y)$ and $R_{q_{2}}(x^{\prime},y)$.
@@
Let $\Delta$ be a countable distance value set with $\inf\Delta=0$. We define a countable $\mathcal{L}$-structure $\mathcal{M}_{\Delta}$. Let $\mathbb{R}^{+}$ denote the semigroup of positive real numbers. Let $\bar{\Delta}\subseteq\mathbb{R}^{+}$ be the sub-semigroup of $\mathbb{R}^{+}$ generated by $\Delta$. Then either $\Delta=\bar{\Delta}$ or $\Delta=\bar{\Delta}\cap(0,{\rm sup}\Delta]$. Let $|\mathcal{M}_{\Delta}|=\{0\}\cup\bar{\Delta}$. Define $0,\leq,+$ in the obvious way. For $q\in\mathbb{Q}^{+}$ and $x,y\in|\mathcal{M}_{\Delta}|$, if either $x=0$ or $y=0$, then $\lnot R_{q}(x,y)$; otherwise, define $R_{q}(x,y)$ iff $q<x/y$. Finally, if $\sup\Delta=+\infty$, set $c=0$; otherwise, set $c={\rm sup}\Delta$. Now it is easy to verify that $\mathcal{M}_{\Delta}$ is a model of $T$. Note that clause (7) holds since $\Delta$ is dense in $(0,\sup\Delta)$. It is also clear that the map $\Delta\mapsto\mathcal{M}_{\Delta}$ gives a Borel reduction from $\sim_{\mathsf{CDVS}_{0}}$ to $\cong_{T}$.
@@
$q<f(x+y)$ $\iff R_{q}(x+y,x^{\prime})$
$\iff\exists q_{1},q_{2}\ q_{1}+q_{2}=q\land R_{q_{1}}(x,x^{\prime})\land R_{q_{2}}(y,x^{\prime})$
@@
Finally, we verify that the map $\mathcal{M}\mapsto\Delta_{\mathcal{M}}$ is a Borel reduction from $\cong_{T}$ to $\sim_{\mathsf{CDVS}_{0}}$. It is clear that the map is Borel. For the reduction, we first claim that for nonzero $x,y\in\mathcal{M}$, $R_{q}(x,y)$ iff $q<f(x)/f(y)$. Suppose $R_{q}(x,y)$. There is $q^{\prime}>q$ with $R_{q^{\prime}}(x,y)$. Let $p<q^{\prime}f(y)$. Then $p/q^{\prime}<f(y)$ and thus $R_{p/q^{\prime}}(y,y_{0})$. Since $R_{q^{\prime}}(x,y)$, by clause (4), we have $R_{p}(x,y_{0})$ and thus $p<f(x)$. This shows $q^{\prime}\leq f(x)/f(y)$ and thus $q<f(x)/f(y)$. Similarly we can prove that if $\lnot R_{q}(x,y)$, then $q\geq f(x)/f(y)$. This finishes the proof of the claim.
@@
This means $\tilde{g}(x)/\tilde{g}(y)=x/y$ and thus there is a positive real number $\alpha$ such that $\tilde{g}(x)=\alpha x$ for all $x\in f_{1}(\mathcal{M}_{1})$. Note that $f_{2}(c^{\mathcal{M}_{2}})=\tilde{g}(f_{1}(c^{\mathcal{M}_{1}}))=\alpha f_{1}(c^{\mathcal{M}_{1}})$. Thus either
$\Delta_{\mathcal{M}_{2}}=f_{2}(\mathcal{M}_{2})=\alpha\cdot f_{1}(\mathcal{M}_{1})=\alpha\cdot\Delta_{\mathcal{M}_{1}}$
@@
$\begin{array}[]{rcl}\Delta_{\mathcal{M}_{2}}&=&f_{2}(\mathcal{M}_{2})\cap(0,f_{2}(c^{\mathcal{M}_{2}})]\\ &=&\alpha\cdot\left\{f_{1}(\mathcal{M}_{1})\cap(0,f_{1}(c^{\mathcal{M}_{1}})]\right\}=\alpha\cdot\Delta_{\mathcal{M}_{1}}.\end{array}$
In either case, $\Delta_{\mathcal{M}_{1}}\sim\Delta_{\mathcal{M}_{2}}$. Conversely, suppose there is a positive real number $\alpha$ such that $\Delta_{\mathcal{M}_{2}}=\alpha\cdot\Delta_{\mathcal{M}_{1}}$. Then $f_{2}(\mathcal{M}_{2})=\alpha\cdot f_{1}(\mathcal{M}_{1})$ and $f_{2}(c^{\mathcal{M}_{2}})=\alpha f_{1}(c^{\mathcal{M}_{1}})$. Let $g=f_{2}^{-1}\alpha f_{1}\cup\{(0,0)\}$. Then a similar argument shows that $g$ is an isomorphism from $\mathcal{M}_{1}$ to $\mathcal{M}_{2}$. ∎
7. Concluding Remarks and Further Questions
We do not know the exact complexity of any of the equivalence relations considered in Theorem 6.2 in the Borel reducibility hierarchy. About $\approx_{\mathsf{CDVS}}$ we have the following observation.
@@ 7. Concluding Remarks and Further Questions
$\approx_{\mathsf{CDVS}}$ is Borel reducible to an isomorphism relation.

### v2

@@ 5.3. The complexity of $\approx_{\mathsf{CDVS}_{0}}$
For any $A\subseteq\mathbb{R}^{\mathbb{N}}$, let $\sim_{A}$ denote $\sim\upharpoonright(A\times A)$.
Thus Theorem 5.6 can be restated as $\sim_{\mathsf{CDVS}_{0}}\,=\ \approx_{\mathsf{CDVS}_{0}}$.
@@
To obtain some upper and lower bounds for $\sim_{\mathsf{CDVS}_{0}}$ in the Borel reducibility hierarchy, we consider another equivalence relation defined and studied by Calderoni–Marker–Motto Ros–Shani [2]. Let $\mathsf{ArGp}$ be the space of all countable Archimedean ordered groups, and let $\cong_{\mathsf{ArGp}}$ denote the isomorphism relation on $\mathsf{ArGp}$. It was proved in [2] that
$^{+}\ <_{B}\ \cong_{\mathsf{ArGp}}\ \leq_{B}\ \cong^{*}_{3,1}.$
@@
To see that $\Delta$ is a Borel reduction witnessing $\cong_{\mathsf{ArGp}}\ \leq_{B}\ \sim_{\mathsf{CDVS}_{0}}$, we use a result of Hion (see Lemma 2.2 of [2]) which states that for any $\in\mathsf{ArGp}, GHG\cong H$ if and only if there is a real number $r$ such that $r\cdot G=H$. It is then clear that $G\cong H$ if and only if there is a positive real number $r$ such that $r\cdot G=H$, if and only if there is a positive real number $r$ such that $r\cdot\Delta(G)=\Delta(H)$, and if and only if $\Delta(G)\sim\Delta(H)$. ∎
Theorem 5.8.
@@
$\mathcal{L}=\{0,c,\leq,+\}\cup\{R_{q}\colon q\in\mathbb{Q}^{+}\}$
@@
  • ++ is a binary function symbol;
@@
In a countable distance value set $\Delta$ with $\inf\Delta=0$, we interpret $c$ to be 0 if $\sup\Delta=+\infty$ and to be $\sup\Delta$ otherwise. For each $q\in\mathbb{Q}^{+}$, $R_{q}$ is interpreted as $R_{q}(x,y)$ if and only if $q<x/y$. Under these interpretations, $\Delta\cup\{0\}$ becomes a countable $\mathcal{L}$-structure.
@@
    $\leq$ is a linear ordering of $|\mathcal{M}|$, 0 is the $\leq$-least element, and for all $^{\prime}\in|\mathcal{M}|\setminus\{0\}, xxx\leq x^{\prime}$ if and only if $x=x^{\prime}$ or $R_{1}(x^{\prime},x)$.
@@
    For all $x,x^{\prime},y\in|\mathcal{M}|\setminus\{0\}$ and $q\in\mathbb{Q}^{+}$, $R_{q}(x+x^{\prime},y)$ if and only if there are $q_{1},q_{2}\in\mathbb{Q}^{+}$ with $q_{1}+q_{2}=q$ such that $R_{q_{1}}(x,y)$ and $R_{q_{2}}(x^{\prime},y)$.
@@
Let $\Delta$ be a countable distance value set with $\inf\Delta=0$. We define a countable $\mathcal{L}$-structure $\mathcal{M}_{\Delta}$. Let $\mathbb{R}^{+}$ denote the semigroup of positive real numbers. Let $\bar{\Delta}\subseteq\mathbb{R}^{+}$ be the sub-semigroup of $\mathbb{R}^{+}$ generated by $\Delta$. Then either $\Delta=\bar{\Delta}$ or $\Delta=\bar{\Delta}\cap(0,{\rm sup}\Delta]$. Let $|\mathcal{M}_{\Delta}|=\{0\}\cup\bar{\Delta}$. Define $0,\leq,+$ in the obvious way. For $q\in\mathbb{Q}^{+}$ and $x,y\in|\mathcal{M}_{\Delta}|$, if either $x=0$ or $y=0$, then $\lnot R_{q}(x,y)$; otherwise, define $R_{q}(x,y)$ if and only if $q<x/y$. Finally, if $\sup\Delta=+\infty$, set $c=0$; otherwise, set $c={\rm sup}\Delta$. Now it is easy to verify that $\mathcal{M}_{\Delta}$ is a model of $T$. Note that clause (7) holds since $\Delta$ is dense in $(0,\sup\Delta)$. It is also clear that the map $\Delta\mapsto\mathcal{M}_{\Delta}$ gives a Borel reduction from $\sim_{\mathsf{CDVS}_{0}}$ to $\cong_{T}$.
@@
$q<f(x+y)$ $\iff R_{q}(x+y,y_{0})$
$\iff\exists q_{1},q_{2}\ q_{1}+q_{2}=q\land R_{q_{1}}(x,y_{0})\land R_{q_{2}}(y,y_{0})$
@@
Finally, we verify that the map $\mathcal{M}\mapsto\Delta_{\mathcal{M}}$ is a Borel reduction from $\cong_{T}$ to $\sim_{\mathsf{CDVS}_{0}}$. It is clear that the map is Borel. For the reduction, we first claim that for nonzero $x,y\in\mathcal{M}$, $R_{q}(x,y)$ if and only if $q<f(x)/f(y)$. Suppose $R_{q}(x,y)$. There is $q^{\prime}>q$ with $R_{q^{\prime}}(x,y)$. Let $p<q^{\prime}f(y)$. Then $p/q^{\prime}<f(y)$ and thus $R_{p/q^{\prime}}(y,y_{0})$. Since $R_{q^{\prime}}(x,y)$, by clause (4), we have $R_{p}(x,y_{0})$ and thus $p<f(x)$. This shows $q^{\prime}\leq f(x)/f(y)$ and thus $q<f(x)/f(y)$. Similarly we can prove that if $\lnot R_{q}(x,y)$, then $q\geq f(x)/f(y)$. This finishes the proof of the claim.
@@
This means $\tilde{g}(x)/\tilde{g}(y)=x/y$ and thus there is a positive real number $r$ such that $\tilde{g}(x)=rx$ for all $x\in f_{1}(\mathcal{M}_{1})$. Note that $f_{2}(c^{\mathcal{M}_{2}})=\tilde{g}(f_{1}(c^{\mathcal{M}_{1}}))=rf_{1}(c^{\mathcal{M}_{1}})$. Thus either
$\Delta_{\mathcal{M}_{2}}=f_{2}(\mathcal{M}_{2})=r\cdot f_{1}(\mathcal{M}_{1})=r\cdot\Delta_{\mathcal{M}_{1}}$
@@
$\begin{array}[]{rcl}\Delta_{\mathcal{M}_{2}}&=&f_{2}(\mathcal{M}_{2})\cap(0,f_{2}(c^{\mathcal{M}_{2}})]\\ &=&r\cdot\left\{f_{1}(\mathcal{M}_{1})\cap(0,f_{1}(c^{\mathcal{M}_{1}})]\right\}=r\cdot\Delta_{\mathcal{M}_{1}}.\end{array}$
In either case, $\Delta_{\mathcal{M}_{1}}\sim\Delta_{\mathcal{M}_{2}}$. Conversely, suppose there is a positive real number $r$ such that $\Delta_{\mathcal{M}_{2}}=r\cdot\Delta_{\mathcal{M}_{1}}$. Then $f_{2}(\mathcal{M}_{2})=r\cdot f_{1}(\mathcal{M}_{1})$ and $f_{2}(c^{\mathcal{M}_{2}})=rf_{1}(c^{\mathcal{M}_{1}})$. Let $g=f_{2}^{-1}rf_{1}\cup\{(0,0)\}$. Then a similar argument shows that $g$ is an isomorphism from $\mathcal{M}_{1}$ to $\mathcal{M}_{2}$. ∎
Corollary 5.9.
$^{+}\ <_{B}\ \approx_{\mathsf{CDVS}_{0}}\ <_{B}\ =^{++}$.
Proof.
Combining with the results of [2] and Theorem 5.6, Lemma 5.7 implies $^{+}\ <_{B}\ \approx_{\mathsf{CDVS}_{0}}$ and Theorem 5.8 implies $\approx_{\mathsf{CDVS}_{0}}\ <_{B}\ =^{++}$. ∎
We also consider a special kind of countable distance value sets for which the equivalence is more explicit.
Definition 5.4.
Let $\alpha$ be a positve irrational number. Let
$\Delta_{\alpha}=\{p\alpha+q:p,q\in\mathbb{Q}\}\cap(0,+\infty).$
Define an action of $GL_{2}(\mathbb{Q})$ on the space of irrational numbers by the linear fractional transformation:
Theorem 5.10.
  1. (1)
    $\Delta_{\alpha}$ and $\Delta_{\beta}$ are equivalent.
  2. (2)
    $\alpha$ and $\beta$ are $E_{GL_{2}(\mathbb{Q})}^{IR}$-equivalent.
Proof.
We note the following fact.
@@
$E_{0}\leq_{B}E_{GL_{2}(\mathbb{Q})}^{IR}\leq_{B}E_{\infty}$.
@@
$E_{GL_{2}(\mathbb{Q})}^{IR}$ is induced by a continuous action of a countable group on a Polish space. Thus $E_{GL_{2}(\mathbb{Q})}^{IR}\leq_{B}E_{\infty}$. Note that every orbit is dense for this action. It follows from a theorem of Becker–Kechris (see Theorem 6.2.1 of [8]) that $E_{0}\leq_{B}E_{GL_{2}(\mathbb{Q})}^{IR}$. ∎
6. Isomorphism of Extremely Amenable Groups in the Framework of Descriptive Set Theory